OpenStudy (anonymous):

Find equations of the tangent lines to the curve y= (x-1)/ (x+1) that are parallel to the line x − 2y = 3.

OpenStudy (anonymous):

I got x=0 and x=-2.

OpenStudy (anonymous):

hi. are you still there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

sorry, inadvertently pressed the wrong button and then i lost u

OpenStudy (anonymous):

oh , ok I got x= 1/2 x- 1 and 1/2 x + 4 but i dont know if it is right or not

OpenStudy (anonymous):

i mean y=*

OpenStudy (anonymous):

i am thinking for a sec

OpenStudy (anonymous):

so for the derivative you got \[y \prime=\frac{ 2 }{ (x+1)^2 }\]?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

..and the slope of the original line is 1/2.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what were your points of tangency for m=1/2?

OpenStudy (anonymous):

what were the x values?

OpenStudy (anonymous):

x= 0 x=-2

OpenStudy (anonymous):

so you set 2/(x+1)^2=1/2 and solved this equation?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

and then i plug in the x values into the original equation and get my y values

OpenStudy (anonymous):

and then i know x,y, and slope, so i just solved for b

OpenStudy (anonymous):

give me a min.

OpenStudy (anonymous):

ok thks

OpenStudy (anonymous):

x=0,-2 is what i got as well.

OpenStudy (anonymous):

wait a sec. those are the y values

OpenStudy (anonymous):

the x values are x=-,1

OpenStudy (anonymous):

u mean 0 and -2 are y values..?

OpenStudy (anonymous):

well, you plugged in x values of -3 and 1 to get them

OpenStudy (anonymous):

in the original equation above

OpenStudy (anonymous):

I set 2/ (x+1)^2 = 2 then (x+1)^2 = 1 x+1=1 and x+1=-1

OpenStudy (anonymous):

no i didnt. my x values are 0 and -2. i didnt plug in anything to get it . i just derived them from the equation

OpenStudy (anonymous):

2/(x+1)^2=1/2

OpenStudy (anonymous):

oh..opps..

OpenStudy (anonymous):

and x= 1 and -3

OpenStudy (anonymous):

x-2y=3->x-3=2y

OpenStudy (anonymous):

ok, you look like you are headed in the right direction

OpenStudy (anonymous):

so the slope we are interested in =1/2

OpenStudy (anonymous):

(1,0) and (-3,2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

nice. you are right there

OpenStudy (anonymous):

now just plug that puppy into the point slope form and you are done

OpenStudy (anonymous):

ok thanks!!! omg, helped me so much.. I dont get math at all..

OpenStudy (anonymous):

yes you do you are very smart. dont think otherwise

OpenStudy (anonymous):

you take care. hope i can help you again sometime

OpenStudy (anonymous):

ok thx!